For vectors A and B (A+B).(A×B) will be _______
Options
A. A2B2
B. (A+B) (AB) 
C. Zero
D. √A2+B2+ AB
 

Grade:11

1 Answers

Arun
4 years ago
Dear student
 
if you open this,
you get A. (A x B ) + B. ( A x B) = 0 + 0 = 0
 
Hope it helps
 
Regards
Arun (askIITians forum expert)